Circular No. 9086

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 U3 (HILL) R. E. Hill reports his discovery of a comet (discovery observation tabu-

lated below) on CCD images taken in 3′′–4′′ seeing with the Catalina 0.68m Schmidt telescope, noting the object in co-added exposures to have an nuclear condensation of diameter ∼ 5′′–7′′ elongated southwest-northeast, with a coma of diameter ∼ 12′′–15′′, and a diffuse tail ∼ 15′′–20′′ long in p.a. 220^{o}–230^{o}. A similar appearance was noted by Hill on four co-added 60-s exposures from Oct. 22.4 UT in slightly better seeing, noting a nuclear condensation of diameter ∼ 10′′–15′′ and “no noticeable coma”, with the broad, diffuse tail being ∼ 20′′ long in p.a. 220^{o}. After posting on the ‘NEOCP’ webpage, other CCD astrometrists have also commented on the object’s cometary appearance. G. Sostero, E. Guido, and P. Camilleri (fifteen co-added unfiltered 60-s exposures obtained remotely with a 0.35-m reflector at the Skylive-Grove Creek Observatory near Trunkey, Australia; Oct. 21.7) report a diffuse coma ∼ 10′′ in diameter with a central condensation and a tail ∼ 12′′ long toward the west; their further co-added frames on Oct. 22.6 also show a diffuse coma of size ∼ 12′′. Exposures taken by A. Asami and N. Hashimoto (Bisei Spaceguard Center 1.0-m reflector; Oct. 21.74–21.76) show a 15′′ coma with a tail 25′′ long in p.a. 235^{o}. P. R. Holvorcem notes that three co-added 120-s exposures taken by M. Schwartz and himself on Oct. 22.4 with the Tenagra II 0.81-m telescope near Nogales, AZ, USA, show a coma diameter of 9′′ with a 5′′ tail toward p.a. 195^{o}. A. D. Grauer (Mt. Lemmon 1.5-m reflector, Oct. 22.4) reported that his coadded 60-s exposures show a 9′′ coma and a flared tail 15′′ long in p.a. 230^{o}. H. Sato (Tokyo, Japan, remotely using a 0.25-m reflector at the RAS Observatory near Mayhill, NM, U.S.A., on Oct. 22.5) finds a diffuse coma of diameter 15′′. R. Miles writes that four stacked 15-s frames taken by G. Privett with the 2.0-m ‘Faulkes Telescope North’ at Haleakala on Oct. 22.6 show the comet to be elongated 5′′.0 × 1′′.5, aligned in p.a. 75^{o}-255^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
Oct. 21.32407 3^{h}50^{m}08.^{s}41 +14^{o}50′23′′.7 17.6 Hill

Additional astrometry, the following parabolic orbital elements, and an ephemeris appear on MPEC 2009-U67:

T = 2010 Mar. 11.739 TT  ω = 83.^{o}676
Ω = 48.951 }2000.0
q = 1.22986 AU  i = 49.134
2009 October 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green